# Foundation Models on a Budget: Approximating Blocks in Large Vision Models

### v5

@@ 4.1 Latent analysis
This suggests that approximating the final blocks would lead to significant changes in the final representations, indicating their critical role. However, in the case of DEiT-S, the trend is different. Here, the MSE decreases toward the final blocks. This is also shown in the similarity matrices illustrated in Figure 2, where the final blocks have a high similarity. These findings reinforce the intuition behind the use of MSE, demonstrating a correlation between the metric and the final representation similarity when approximating blocks. In some instances, such as with the MNIST dataset, the MSE scores remain relatively consistent across blocks, indicating that the representations are largely similar to one another. However, for more complex datasets like ImageNet1k, the representations in the final or the first blocks become increasingly dissimilar, making it advantageous to approximate intermediate blocks. This suggests that the similarities are primarily driven by the model but also guided by the complexity of the dataset, allowing for targeted approximations that reduce model parameters and complexity without significantly compromising performance. To further investigate the effect of TBA on the latent representations, we visualize, using Principal Component Analysis (PCA), the representations of the TBA model when approximating the final block and its original representation (see Figure 4). The plot shows the representations generated using the DiNO-B and DEiT-S pretrained encoders on F-MNIST. Colors represent classes. The visualization highlights that approximating the final block of DiNO-B results in noticeable deviations from the original representations, while in DEiT-S the approximated representations remain similar to the original ones. This observation aligns with the results from Figure 3, where approximating blocks can lead to significant changes in representations. For complete results, refer to Section A.2.2.
4.2 Transformer block approximation performance
We perform image classification using various pretrained models with different dimensionalities (ViT-T, ViT-S, ViT-L, ViT-B, DiNO-S, DiNO-B, and DEiT-S) and datasets (MNIST, F-MNIST, CIFAR-10, CIFAR-100C, CIFAR-100F, and ImageNet1k), keeping all models frozen. Block approximations are computed using a shared linear transformation applied across all tokens, based on
| Accuracy | |||||
| Approx. | Params. | CIFAR-10 | CIFAR-100F | ImageNet1k | |
| ViT-S | 1 5 | -6.51M | (-11.09) | (-27.35) | (-39.86) |
| 2 5 | -4.88M | (-4.88) | (-14.42) | (-17.56) | |
| 7 10 | -4.88M | (-8.69) | (-23.89) | (-51.14) | |
| 2 4 | -3.26M | (-2.95) | (-8.67) | (-6.80) | |
| 9 11 | -3.26M | (-5.00) | (-14.71) | (-35.54) | |
| 1 2, 4 5 | -3.26M | (-3.54) | (-11.10) | (-12.06) | |
| 0 1 | -1.63M | (-2.09) | (-6.15) | (-11.10) | |
| 3 4 | -1.63M | (-0.66) | (-2.47) | (-2.51) | |
| 9 10 | -1.63M | (-1.15) | (-6.19) | (-15.58) | |
| 10 11 | -1.63M | (-1.67) | (-4.97) | (-12.73) | |
| original | 21.82M | ||||
| DiNO-B | 1 5 | -26.5M | (-11.28) | (-28.62) | (-58.57) |
| 2 5 | -19.5M | (-5.87) | (-19.50) | (-28.26) | |
| 7 10 | -19.5M | (-4.46) | (-14.02) | (-35.88) | |
| 2 4 | -13M | (-1.49) | (-6.62) | (-7.70) | |
| 9 11 | -13M | (-10.76) | (-14.36) | (-27.74) | |
| 1 2, 4 5 | -13M | (-1.47) | (-6.11) | (-7.79) | |
| 0 1 | -6.5M | (-0.05) | (-1.01) | (-1.00) | |
| 3 4 | -6.5M | (-0.38) | (-2.78) | (-1.62) | |
| 9 10 | -6.5M | (-0.78) | (-3.26) | (-10.09) | |
| 10 11 | -6.5M | (-1.01) | (-4.23) | (-13.83) | |
| original | 86.58M | ||||
| DEiT-S | 1 5 | -6.51M | (-13.55) | (-28.90) | (-40.88) |
| 2 5 | -4.88M | (-5.41) | (-14.87) | (-16.05) | |
| 7 10 | -4.88M | (-1.61) | (-2.78) | (-22.24) | |
| 2 4 | -3.26M | (-1.85) | (-6.37) | (-5.32) | |
| 9 11 | -3.26M | (+0.30) | (+1.11) | (-5.39) | |
| 1 2, 4 5 | -3.26M | (-5.74) | (-13.31) | (-10.68) | |
| 0 1 | -1.63M | (-6.21) | (-12.91) | (-15.38) | |
| 3 4 | -1.63M | (-0.14) | (-1.24) | (-1.22) | |
| 9 10 | -1.63M | (+0.30) | (+0.86) | (-2.15) | |
| 10 11 | -1.63M | (+0.49) | (+1.15) | (+0.05) | |
| original | 21.82M | ||||
a subset of 3,000 training samples. Then, we train a single linear layer using the Adam optimizer (learning rate ) for 5 epochs, with 3 seeds and a batch size of 256. Details regarding the models, datasets, computational resources, and software tools are provided in Table 6, Table 7, Section A.1.4, and Section A.1.3, respectively.
Image classification As presented in Table 1, the proposed TBA method successfully reduces model size while maintaining, and in some cases even improving, image classification performance. These downstream task results support the findings from the analysis in Section 4.1. Specifically, as discussed in Section 4.1 and illustrated in Figures 3 and 4, approximating the final block of DEiT-S yields a final representation highly similar to the original. This similarity suggests such blocks are optimal candidates for approximation, a conclusion further validated by Table 1. The table shows that even when approximating multiple consecutive blocks (e.g., 911), we can reduce parameters while achieving performance superior to the original model. Overall, the consistent or enhanced performance indicates that a linear transformation is sufficient to approximate these transformer blocks, thereby effectively reducing model parameters. It’s important to note that this transformation is shared across all tokens, with no additional training steps required. Additional results are presented in Tables 8, 9, 10, 11, 12 and 13. To assess the scalability of our method, we extended our evaluation to ViT-L. While our approach reduces its parameters, ViT-L still imposes considerable computational demands, necessitating high-performance GPU infrastructure. This finding underscores that while effective for substantial compression, our method’s impact on making extremely large models practical for resource-constrained environments is limited by their initial scale. Finally, in Section A.2.6, we further analyze model behavior after block approximations on the final image-classification task, complementing the quantitative results with qualitative insights into the approximation procedure.
Zero-shot image classification To further assess the effectiveness of our approach, we evaluate
| Params. | Approx. | Accuracy | |
| -6.49M | 0 1 | 57.93 | - |
| 1 2 | 64.20 | - | |
| 2 3 | 66.35 | -5.51% | |
| 3 4 | 64.65 | - | |
| 4 5 | 64.86 | - | |
| 5 6 | 58.05 | - | |
| 6 7 | 61.56 | - | |
| 7 8 | 58.53 | - | |
| 8 9 | 52.32 | - | |
| 9 10 | 59.21 | - | |
| 10 11 | 22.64 | - | |
| 149.07M | original | 70.21 | – |
TBA in a zero-shot image classification setting. This evaluation utilizes the OpenCLIP-ViT-B model (Radford et al., 2021), which was pretrained on LAION-2B Schuhmann et al. (2022), with ImageNet1k serving as the downstream evaluation dataset. As in previous experiments, the model remains frozen, and block approximations are computed using a shared linear transformation applied across all tokens, based on a subset of 3,000 training samples. Importantly, we apply these approximations only to the vision encoder, leaving the text encoder unchanged. We follow the standard ImageNet1k prompt templates. As shown in Table 2, when approximating earlier vision blocks (e.g., 1 2 or 2 3), the model still achieves competitive results, but slightly reduces the number of parameters. The analysis is conducted only on the base version, as larger versions (e.g., OpenCLIP-ViT-L or OpenCLIP-ViT-H) contain too many parameters and are thus beyond the scope of this paper. Details on the models are provided in Table 6.
Transformation generalization We further evaluate the model’s ability to generalize transformations across datasets. Specifically, we analyze the capability of applying a linear transformation learned on one dataset (source) to another (target), using the same underlying pretrained architecture and without any subsequent training or fine-tuning of the approximated model. Table 3 reports accuracy scores for ViT-S and DiNO-S on MNIST, CIFAR-10, and CIFAR-100, where the "Source" column indicates the source dataset used to compute the transformation. The results demonstrate that a single, token-shared linear transformation can generalize effectively across different target datasets, with the exception of MNIST, which may be too simplistic to yield broadly transferable representations. This cross-dataset generalization is particularly valuable in scenarios where fine-tuning is infeasible, such as in privacy-sensitive domains like healthcare or finance. In these contexts, regulatory constraints often restrict training on local data, yet leveraging existing pretrained models remains crucial. Our findings suggest that lightweight adaptation via shared linear transformations offers a promising path for improving model performance on custom datasets without requiring additional training. Additional results are provided in Tables 14 and 15.
| Accuracy | ||||
| Approx. | Source | MNIST | CIFAR-10 | CIFAR-100C |
| 3 4 | MNIST | |||
| CIFAR-10 | ||||
| CIFAR-100F | ||||
| 3 5 | MNIST | |||
| CIFAR-10 | ||||
| CIFAR-100F | ||||
4.3 Comparison
Direct comparisons for TBA are limited, as existing approaches either fine-tune pretrained models or train from scratch. In contrast, TBA maintains the foundation model frozen. We include two types of comparisons: a variant of SkipAt, a method proposed in Venkataramanan et al. (2024), and an ablation study on transformation complexity.
SkipAt identity Venkataramanan et al. (2024) explores the effect of skipping self-attention or multi-head self-attention by inserting identity functions in pretrained models, without retraining. We extend this idea by skipping entire blocks, not just attention components. As shown in Table 4, TBA consistently outperforms this SkipAt Identity baseline on CIFAR-10 and CIFAR-100F. The accuracy improvements confirm that approximating is more effective than skipping them. Results with other datasets and models can be found in Section A.2.5.
| CIFAR-10 | CIFAR-100F | |||
| Approx. | SkipAt | TBA | SkipAt | TBA |
| 1 5 | ||||
| 2 5 | ||||
| 7 10 | ||||
| 1 3 | ||||
| 2 4 | ||||
| 9 11 | ||||
| 2 3 | ||||
| 9 10 | ||||
| original | ||||
Trained approximators We also evaluate whether more complex approximations improve performance. Specifically, we compare TBA to deeper MultiLayer Perceptron (MLP)-based translators (MLP and Res-MLP), which are trained for 300 steps using Adam with a learning rate of 1e-3. As shown in Table 5, employing a simple linear transformation to approximate blocks is the optimal choice in most cases. While deeper translators occasionally yield slightly higher accuracy (e.g., 15), they incur significant training overhead and require gradient-based optimization. In contrast, TBA is entirely training-free and operates in closed form; no backpropagation or updates are required. This highlights the efficiency–performance trade-off and supports the design choice behind TBA. Refer to Section A.1.2 for approximator details.
| Approx. | TBA | MLP | Res-MLP |
| 1 5 | |||
| 2 5 | |||
| 7 10 | |||
| 1 3 | |||
| 2 4 | |||
| 9 11 | |||
| 3 4 | |||
| 9 10 |
@@ 5 Limitations and future work
While TBA demonstrates significant promise in efficiently approximating transformer blocks, our current investigation has primarily focused on vision transformer architectures and their application to classification tasks. Future research will explore the applicability of TBA to other modalities (e.g., text) and to diverse downstream tasks (e.g., image reconstruction). Such an expansion will be crucial for testing the universality of the observed block-similarity phenomena and assessing TBA’s adaptability. Furthermore, we aim to expand the analysis of these block-level similarities. This involves investigating redundancies at finer granularities, such as within individual attention heads or feed-forward network layers, and potentially at coarser granularities across larger segments of the network. Another important direction is to examine more in detail how these similarity patterns arise depending on token type (e.g., CLS token versus last token). The analysis could lead to even more refined and context-aware approximation strategies, further enhancing model efficiency.

### v6

@@ 4.1 Latent analysis
To quantify the effect of the approximation, we compute the CKA similarity between the final block representations of the original and the TOAST-approximated model for each block using its preceding block as input. As shown in Figure˜4, the model-specific similarity patterns re-emerge after approximation. The plots highlight more specific trends. Approximating blocks is easier on simpler tasks (e.g., image classification on MNIST or F-MNIST), yielding representations that closely match the originals, whereas on more complex datasets (e.g., ImageNet1k or CIFAR-100), the approximated representations deviate more from the original ones. Furthermore, the final blocks of DEiT-S exhibit high similarity, suggesting that approximating these layers preserves the final representations, while earlier blocks remain more critical.
To provide a more intuitive view, Figure˜4 visualizes the final-layer representations using Principal Component Analysis (PCA). We compare the original representations with those obtained after approximating the final block (10 11) using TOAST on F-MNIST, with colors indicating the 10 classes. The visualization confirms that approximating the final block of DiNO-B results in noticeable deviations from the original representations, whereas for DEiT-S the approximated representations remain highly similar. These observations align with the CKA analysis in Figure˜4, highlighting that the effect of block approximation depends strongly on the model and its internal block structure. Additional results across other models and datasets are provided in Section˜A.2.1.
| Identity Translator | Linear Translator | |||||||
| Approx. | Params. | Accuracy % | GFLOPS | imgs/s | Accuracy % | GFLOPS | imgs/s | |
| 2 4 | -3.25M | |||||||
| 9 11 | -3.25M | |||||||
| 0 1 | -1.62M | |||||||
| 10 11 | -1.62M | |||||||
| original | 21.81M | 4.97 | 6349.2 | 4.97 | 6325.6 | |||
Can entire transformer blocks be approximated without losing accuracy?
Initial results, reported in Table˜1, support the qualitative analysis and empirically demonstrate that entire vision transformer blocks can be effectively approximated using simpler transformations (e.g., linear projections or, in some cases, the identity function). Such approximations reduce both the number of parameters and Giga Floating-Point Operations (GFLOPs), thereby improving throughput (images per second), while incurring only a slight to negligible decrease in downstream task performance. For instance, consistent with our earlier analysis, we find that approximating the final block of DEiT-S when using ImageNet1k (e.g., approximating blocks 10 11 or 9 11 with a linear transformation) yields modest performance drops of -0.09% and -3.38%, respectively, while providing substantial efficiency gains. Importantly, we also show that even the identity transformation, achieves competitive results, with accuracy drops as small as -0.24% and -5.44%, respectively. However, the choice of translator naturally depends on the efficiency-accuracy trade-off: linear translation guarantee in general most reliable accuracy–efficiency balance, whereas the identity yields the leanest training-free approximation when maximum simplicity is required. Further methodological details and the full evaluation are presented in Section˜A.1 and Section˜4.2 respectively, while details on the efficiency metrics and additional analysis on those are in Sections˜A.1.5 and A.2.5, respectively.
4.2 Image Classification Performance
We evaluate TOAST on image classification tasks using pretrained models of varying sizes (ViT-L, DiNO-B, and DEiT-S) and two benchmark datasets (CIFAR-100F and ImageNet1k). Additional results with a broader set of models (ViT-T, ViT-S, ViT-B, ViT-L, DiNO-S, DiNO-B, DEiT-S) and datasets (MNIST, F-MNIST, CIFAR-10, CIFAR-100C) are provided in Section˜A.2.2. While in Section˜A.2.6, we complement the quantitative evaluations with qualitative analyses of misclassifications after block approximation, providing further insight into model behavior under TOAST. Additional implementation details, including model and dataset specifications, computational resources, and software tools, are provided in Tables˜5 and 6, and Sections˜A.1.4, A.1.5 and A.1.3.
| Identity | Linear | |||||||
| Approx. | Params. | Accuracy % | GFLOPs | imgs/s | Accuracy % | GFLOPs | imgs/s | |
DEiT-S |
3 4, 9 11 | -4.88M | 3.74 | 7751.4 | 3.80 | 7718.4 | ||
| 3 4, 9 10 | -3.25M | 4.15 | 7210.9 | 4.21 | 7188.4 | |||
| 2 3 | -1.62M | 4.56 | 6754.2 | 4.59 | 6736.7 | |||
| 10 11 | -1.62M | 4.56 | 6752.6 | 4.59 | 6740.5 | |||
| original | 21.81M | 4.97 | 6349.2 | 4.97 | 6325.6 | |||
DiNO-B |
0 4 | -26.00M | 16.32 | 3230.9 | 16.47 | 3227.7 | ||
| 0 1, 2 3, 4 5 | -19.50M | 18.34 | 2947.0 | 18.80 | 2929.6 | |||
| 0 1, 2 3 | -13.00M | 20.37 | 2703.9 | 20.67 | 2681.2 | |||
| 0 1 | -6.50M | 22.39 | 2506.6 | 22.54 | 2487.0 | |||
| 5 6 | -6.50M | 22.39 | 2503.1 | 22.54 | 2490.6 | |||
| original | 86.58M | 24.42 | 2321.3 | 24.42 | 2316.5 | |||
ViT-L |
2 4, 18 23 | -80.83M | 45.05 | 1654.9 | 45.47 | 1652.8 | ||
| 17 23 | -69.28M | 47.70 | 1572.4 | 47.90 | 1567.0 | |||
| 3 4, 19 23 | -57.74M | 50.34 | 1509.9 | 50.75 | 1499.5 | |||
| 3 4, 20 23 | -46.19M | 52.98 | 1440.4 | 53.39 | 1436.8 | |||
| 3 4, 21 23 | -34.64M | 55.62 | 1377.2 | 56.03 | 1345.6 | |||
| 7 8, 15 16 | -23.09M | 58.26 | 1318.2 | 58.67 | 1316.4 | |||
| 16 17, 22 23 | -23.09M | 58.26 | 1318.8 | 58.67 | 1312.3 | |||
| 3 4 | -11.55M | 60.90 | 1269.2 | 61.11 | 1270.0 | |||
| 22 23 | -11.55M | 60.90 | 1267.5 | 61.11 | 1270.9 | |||
| original | 304.35M | 63.54 | 1219.8 | 63.54 | 1225.2 | |||
Block approximations in TOAST are calculated via a shared linear, or identity, transformation applied across all tokens and are estimated using a subset of 500 training samples. A linear classifier is then trained on top of the frozen backbone with the Adam optimizer (learning rate ), batch size 256, for 5 epochs, over 3 different seeds. This setup simulates a realistic scenario where a pretrained feature extractor is adapted to a new dataset unseen during pretraining. However, to assess the robustness of our method, we also report the results using the original classification heads (Section˜A.2.4), which confirm the consistency of our findings.
Are TOAST results competitive?
As shown in Tables˜3 and 2, TOAST consistently reduces model size and GFLOPs while maintaining, and in some cases improving, image classification accuracy. This aligns with our representational analyses in Section˜4.1: for instance, approximating the final block of DEiT-S produces latent representations nearly identical to the original (Figures˜4 and 4), making it an ideal candidate for approximation. Even when multiple consecutive blocks are approximated (e.g., 911), models maintain performance comparable to or exceeding the original while significantly reducing parameters. This demonstrates that a simple linear transformation, or even the identity in certain cases, is sufficient to capture the functionality of full transformer blocks without additional training, provided the transformation is shared across all tokens. Additionally, efficiency gains are notable: throughput (imgs/s) increases while GFLOPs decreases, highlighting practical benefits for deployment, as also shown in Section˜A.2.5. Additional results across other models (DiNO-B, ViT-L) and datasets confirm that TOAST generalizes across architectures and scales (Section˜A.2.2). Finally, while approximations are easier for simpler datasets (e.g., CIFAR-100F), TOAST still achieves meaningful compression with minimal accuracy loss on complex datasets like ImageNet1k.
| Identity | Linear | |||||||
| Approx. | Params. | Accuracy % | GFLOPs | imgs/s | Accuracy % | GFLOPs | imgs/s | |
DEiT-S |
3 4, 9 11 | -4.88M | 3.74 | 7755.1 | 3.80 | 7713.7 | ||
| 9 11 | -3.25M | 4.15 | 7226.6 | 4.18 | 6791.7 | |||
| 8 9 | -1.62M | 4.56 | 6755.2 | 4.59 | 6739.9 | |||
| 9 10 | -1.62M | 4.56 | 6692.1 | 4.59 | 6741.3 | |||
| original | 21.81M | 4.97 | 6349.0 | 4.97 | 6249.4 | |||
DiNO-B |
0 4 | -26.00M | 16.32 | 3233.8 | 16.47 | 3204.9 | ||
| 0 1, 2 3, 4 5 | -19.50M | 18.34 | 2943.1 | 18.80 | 2922.6 | |||
| 0 1, 2 3 | -13.00M | 20.37 | 2705.6 | 20.67 | 2690.1 | |||
| 0 1 | -6.50M | 22.39 | 2492.8 | 22.54 | 2493.8 | |||
| 2 3 | -6.50M | 22.39 | 2484.2 | 22.54 | 2484.6 | |||
| original | 86.58M | 24.42 | 2315.5 | 24.42 | 2317.3 | |||
ViT-L |
2 4, 18 23 | -80.83M | 45.05 | 1655.7 | 45.47 | 1649.6 | ||
| 17 23 | -69.28M | 47.69 | 1578.8 | 47.90 | 1552.1 | |||
| 3 4, 19 23 | -57.74M | 50.34 | 1503.6 | 50.75 | 1497.4 | |||
| 3 4, 20 23 | -46.19M | 52.98 | 1445.2 | 53.39 | 1431.0 | |||
| 20 23 | -34.64M | 55.62 | 1381.2 | 55.82 | 1372.6 | |||
| 3 4, 21 23 | -34.64M | 55.62 | 1376.7 | 56.03 | 1372.7 | |||
| 20 22 | -23.09M | 58.26 | 1317.5 | 58.47 | 1314.6 | |||
| 3 4, 21 22 | -23.09M | 58.26 | 1315.8 | 58.67 | 1317.6 | |||
| 20 21 | -11.55M | 60.90 | 1268.5 | 61.11 | 1266.7 | |||
| 21 22 | -11.55M | 60.90 | 1270.7 | 61.11 | 1269.2 | |||
| original | 304.35M | 63.54 | 1223.1 | 63.54 | 1224.3 | |||
Additional results across models and datasets are provided in Tables˜7, 8, 9 and 10. To assess scalability, we applied TOAST to ViT-L. Approximating selected blocks, e.g., 17 23, reduces the parameter count by 69.3M, lowers GFLOPs from 63.54 to 47.79, and increases throughput from 1223.1 to 1578.8 imgs/s, while incurring a minimal accuracy drop of 1.16%. This shows TOAST’s utility in balancing substantial computational savings with a modest performance trade-off, even in large models.
Are 500 training samples enough?
We study the sensitivity of block approximation to the number of training samples using DiNO-B and DEiT-S on ImageNet1k. As shown in Figure˜5, performance typically plateaus quickly: 500 samples are sufficient to obtain stable and reliable approximations. Increasing the sample count beyond this threshold provides only marginal gains, while substantially fewer samples lead to noticeable degradation. Interestingly, when the representational spaces of consecutive blocks are already highly aligned, even as few as 10 or 50 samples suffice to achieve competitive approximations. Conversely, for blocks that are harder to approximate, such as the early layers of DEiT-S (e.g., 01), even 4000 samples are insufficient to estimate a linear transformation that maintains competitive performance. We highlight that these results are obtained on ImageNet1k, which contains 1000 classes. The 500 samples represent only a small subset of the class space, yet reliable approximations are still achieved. This indicates that TOAST primarily captures the block-level structure of representations rather than requiring exhaustive coverage of all classes. Consequently, TOAST could be practical also in scenarios where a large labeled datasets is limited.
What if a more complex transformation is used?
We evaluate whether deeper approximators improve downstream task performance. Specifically, we compare TOAST (Identity and Linear) to MultiLayer Perceptron (MLP) and Residual MLP, trained for 300 steps with Adam (learning rate ). These more complex transformation, as for Identity and Linear, are applied across all tokens, and estimated using a subset of 500 training samples. Results in Table˜4 show a consistent trend for ViT-L on both ImageNet1k and CIFAR-100F: the linear transformation provides the most reliable trade-off across datasets. On CIFAR-100F, linear often achieves the best or near-best accuracy (e.g., 2122: vs. for Res-MLP and for MLP), while remaining training-free, thus more efficient. On ImageNet1k, the gap becomes even clearer: for the same blocks linear reaches , while Res-MLP and MLP reach and , respectively. Additionally, also Linear obtain competitive results. TOAST operates in closed form, requires no optimization, and consistently achieves strong efficiency–accuracy trade-offs. These findings confirm that a simple linear transformation is sufficient to approximate transformer blocks in most settings, with deeper translators offering little benefit despite their higher cost.
| Accuracy | ||||||
| Approx. | Params. | Identity | Linear | MLP | Res-MLP | |
CIFAR-100F |
3 4, 20 23 | -46.19M | ||||
| 17 23 | -69.28M | |||||
| 3 4, 19 23 | -57.74M | |||||
| 20 23 | -34.64M | |||||
| 20 22 | -23.09M | |||||
| 3 4, 21 22 | -23.09M | |||||
| 20 21 | -11.55M | |||||
| 21 22 | -11.55M | |||||
| original | 304.35M | |||||
ImageNet1k |
3 4, 20 23 | -46.19M | ||||
| 17 23 | -69.28M | |||||
| 3 4, 19 23 | -57.74M | |||||
| 20 23 | -34.64M | |||||
| 20 22 | -23.09M | |||||
| 3 4, 21 22 | -23.09M | |||||
| 20 21 | -11.55M | |||||
| 21 22 | -11.55M | |||||
| original | 304.35M | |||||
@@ 5 Limitations and future work
While TOAST efficiently approximates transformer blocks, our current investigation has primarily focused on vision transformer architectures and their application to classification tasks. Future research will explore the applicability of TOAST to other modalities (e.g., text) and to diverse downstream tasks (e.g., image reconstruction). Such an expansion will be crucial for testing the universality of the observed block-similarity phenomena and assessing TOAST’s adaptability. Furthermore, we aim to expand the analysis of these block-level similarities. This involves investigating redundancies at finer granularities, such as within individual attention heads or feed-forward layers, and consistently and automatically identifying which blocks to approximate, resulting in minimal degradation of downstream task performance. Another important direction is to examine more in detail how these similarity patterns arise depending on token type (e.g., CLS token versus mean token). The analysis could lead to even more refined and context-aware approximation strategies, further enhancing model efficiency.